A continuous aligned fibre-reinforced composite is produced containing 36% aramid fibres of tensile strength 2.24GPa and the remainder of polycarbonate with tensile strength 99 MPa.  What is the longitudinal strength of the composite in MPa?


1
Expert's answer
2021-01-11T11:42:43-0500

Answer

fibre-reinforced composite is produced containing 36% aramid fibres of tensile strength 2.24GPa

And

the remainder of polycarbonate with tensile strength 99 MPa.

Crifical fibre length is given by

"l_c=\frac{\sigma d}{2\tau}=\frac{3500\times0.015}{2 \times100 }0.26mm"

Now the longitudinal strength of the composite is given by

"\sigma=\sigma_fV_f(1-\frac{l_c}{2L}) +\sigma_m(1-V_f)"

Using all given data in question

"\sigma=822MPa"